Operations Report – May 5th

Crew 212 Operations Report 05-MAY-2019

SOL: 0

Name of person filing report: Paolo Guardabasso

Non-nominal systems: Loft pump

Notes on non-nominal systems: The loft pump in the Hab is leaking
badly, and water penetrated the wooden floor down in the commander’s
room. The leaking is due to parts of the pump not being completely
fixed to each other and is triggered mainly by vibrations during
activity; water both drops on the wooden floor but also sprays upwards
close to electrical connections.

Generator (hours run last night): 12:20
On (last night): 6:40 PM
Off (this morning): 7:00 AM

Solar— SOC 85% at 7:10 PM

Diesel Reading – 99%

Propane Reading – 49%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 370 gallons

Axillary to Static tank transfer – no

Gallons transferred: 100 gallons

Water in GreenHab – 190 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 1432449

Toilet tank emptied: No

Deimos rover used: no (out of order)
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 92.6h
Beginning charge: 100%
Ending charge: 83%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 58.5h
Beginning charge: 99%
Ending charge: 88%
Currently charging: No

Curiosity rover used: Yes
Hours: 93.8h
Beginning charge: 100%
Ending charge: 90
Currently charging: Yes

Notes on rovers: Opportunity’s battery overheat alarm was on and the
rover stopped functioning on the way back from the short EVA training
with the Director. Opportunity will be moved and plugged for charging
by the Director tonight, before sim begins tomorrow.

ATV’s Used: –

Reason for use: –

Oil Added? –

ATV Fuel Used: –

# Hours the ATVs were Used today: –

Notes on ATVs: –

HabCar used and why, where? A trip to Hanksville was done to collect
100 gallons of water, and another one will be done later tonight with
the same objective.

CrewCar used and why, where? Nothing to report

General notes and comments: Today we had the pre-mission briefing with
the Director

Summary of internet: Nothing to report

Summary of suits and radios: Tested and ready for EVA

Summary of Hab operations: A major cleaning of the Hab was done today,
and also a food inventory

Summary of GreenHab operations: We calculated the correct total volume
of the water tank (250 gallons) and an inventory was prepared

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: The loft pump
probably has to be detached from the main system, disassembled and
checked in case it has to be substituted.

Copyright © The Mars Society. All rights reserved. | Main Site